# Locally rigid $\infty$-categories

### v1

@@ Abstract
We survey the theory of locally rigid and rigid symmetric monoidal $\infty$-categoriesover an arbitrary base $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$. Along the way, we introduce and study “$\mathcal{V}$-atomic morphisms”, which are analogues of compact morphisms over an arbitrary base $\mathcal{V}$.
@@ Introduction
Typically, one studies not only compactly generated $\infty$-categories, but also (symmetric) monoidal versions of them. Among those, the really well behaved ones are the compactly rigidly generated $\infty$-categories, namely those whose compact objects are all dualizable. Typical examples include the $\infty$-category of spectra, or more generally the $\infty$-category of $R$-module spectra for any commutative ring spectrum $R$, but also slightly more exotic objects such as the genuine equivariant stable homotopy category $\mathbf{Sp}_{G}$.
The rise of dualizable $\infty$-categories therefore requires the rise of a notion of symmetric monoidal dualizable category analogous to this one, without requiring a sufficient supply of dualizable objects. This notion is that of a rigid $\infty$-category, as introduced by Gaitsgory and Rozenblyum, see e.g. [GR19a, Chapter 1, §9] or [HSSS21, Section 2.2] for accounts. We view rigid $\infty$-categories as presentably symmetric monoidal $\infty$-categories that are as close as they could be to being “rigidly-compactly generated” without quite being compactly generated. A good example to keep in mind is the $\infty$-category $\mathbf{Sh}(X)$ of sheaves of spectra on a compact Hausdorff space $X$.
The main goal of this paper is to survey in some depth the foundations of the notion of rigid $\infty$-categories, and how they relate to dualizable $\infty$-categories. As in [Ram24] [^1] Of which the present paper can be viewed as a multiplicative companion, we use a presentability result as a pretext for writing about rigid categories, namely, our main result is perhaps the following:
@@
Along the way, we are naturally lead to the study of atomic morphisms, which are to atomic objects what compact morphisms are to compact objects - or maybe they are to compact morphisms what atomic objects are to compact objects (see [Ram24] for the relevant notions of atomic objects and compact morphisms). In any case, they are a relative notion of compact morphisms. A second goal of this paper is also to exposit the basic properties of atomic morphisms, and try to explain how one might manipulate them - they are slightly more finnicky than compact morphisms in the stable setting, though they have the advantage of working in large generality. To illustrate the analogy, we prove the following theorem, which is meant to generalize [Ram24, Theorem 2.39] - see 2.23 for precise definitions:
@@
The precise goal of this paper is therefore to study the theory of locally rigid $\infty$-categories, and relate it to that of rigid $\infty$-categories and atomic morphisms. We also have a presentability result in this case, though because of the non-compactness/atomicity of the unit, one has to be a bit more careful in stating it. Instead, the pretext theorem for many of the developments is the following result which precisely relates locally rigid and rigid $\infty$-categories- it can be viewed as stating that locally rigid $\infty$-categories all arise as “completions” of rigid ones:
@@
In Section 1, we review preliminaries concerning weighted colimits in enriched $\infty$-categories, which will be useful later on to set up the notion of “atomic presentation”, which is our relative analogue of the compact exhaustions of [Ram24, Section 2.2].
In Section 2, we dive into the heart of the matter by introducing atomic morphisms and studying their basic properties, which are meant to be analogous to those of compact morphisms. This is where we prove 2.25, namely B from the introduction.
In Section 3, we recall the definition and basic properties of trace-class morphisms which will be crucial to the study of rigidity.
Section 4 is the main section of the paper, this is where we introduce and study local rigidity and rigidity. We extend and give new proofs of several results from [AGK+20, Appendix C], and give an analogue of 2.25 for rigid categories, namely 4.94. Throughout this section, we explain how the story simplifies if we fix $\mathcal{V}=\mathbf{Sp}$, and prove the relevant analogous results. Finally, it is in this section that we prove the existence of, and give a “concrete” construction of rigidifications, allowing us to prove 4.63, namely C from the introduction.
Finally, in Section 5, we do some set theory and prove 5.4, namely A, as well as its more technical locally rigid analogue, 5.5.
@@
As mentioned in the introduction, the notion of rigid $\infty$-category comes from the work of Gaitsgory and Rozenblyum [GR19a], and was further studied in [HSSS21]; and that of locally rigid $\infty$-category comes from [AGK+20]. Many of the results we prove here can be found there in some form or other, and part of this work can be seen as an organization of these facts, as well as a generalization to arbitrary bases.
Much of it is also inspired by the work of Clausen and Scholze on nuclear categories, and some of it also comes directly from the horse’s mouth, such as 4.56 and the corresponding 4.74, that I learned directly from Dustin Clausen.
@@ Acknowledgements
The acknowledgements for this paper are essentially the same as for the companion [Ram24]. I specifically need to thank Robert Burklund for asking specifically about locally rigid categories rather than restricting to rigid ones.
@@ 1 Preliminaries: weighted colimits
where the first equivalence comes from 1.1.
@@
This allows to define weighted colimits more generally, as objects satisfying this equivalence if they exist [^5] This is beyond the scope of this paper, but this allows to also define “absolute weights”, which are weights $W$ such that $\mathrm{colim}_{I}^{W}(-)$ is preserved whenever it exists. Once this notion is set up, one may try to compare Karoubi complete $\mathcal{V}$-categories, that is, those $\mathcal{V}$-categories that admit all absolute weighted colimits, and atomically generated $\mathcal{V}$-modules. They should be equivalent along the $\mathcal{P}_{\mathcal{V}}\dashv(-)^{\mathrm{at}}$ adjunction from [BM23]..
@@
Let $\underline{\mathbf{1}}:(\mathbf{1}_{\mathcal{W}}\otimes I)^{\mathrm{op}}\to\mathcal{V}$ be the constant weight with value $\mathbf{1}_{\mathcal{V}}$. In this case, combining universal properties we find that $\mathrm{colim}_{\mathbf{1}_{\mathcal{V}}\otimes I}^{\underline{\mathbf{1}}}(-):\operatorname{Fun}(I,\mathcal{M})\simeq\operatorname{Fun}_{\mathcal{V}}(\mathbf{1}_{\mathcal{V}}\otimes I,\mathcal{M})\to\mathcal{M}$ is the ordinary $I$-indexed colimit functor.
@@
If $f$ is strong symmetric monoidal and has a right adjoint $f^{R}$, $f_{\#}$ is left adjoint to $(f^{R})_{\#}$, and we let $f^{*}$ denote the latter.
@@
One can also give a proof at the level of cateories that the resulting functors are $\mathcal{V}$-naturally equivalent.
@@ 2 Atomic maps
In particular, when $\mathcal{V}=\mathbf{Sp}$, we will see that $\mathbf{Sp}$-atomic maps are closely related to compact maps (see 2.6 for a precise statement).
@@
The first half of the claim is 2.5. The second half follows from the first together with [Ram24, Example 2.20]. ∎
@@
As a consequence, ${\mathrm{at}}_{\mathcal{V}}(x,-)\simeq\hom_{\mathcal{V}}(x,\mathbf{1}_{\mathcal{V}})\otimes-$, and so $\mathcal{V}$-atomic maps in $\mathcal{V}$ are the same as trace-class maps (see 3.1 and 3.2).
@@
We can describe ${\mathrm{at}}_{\mathcal{M}}(x,-)$ a bit more explicitly when $\mathcal{M}$ is dualizable over $\mathcal{V}$ in fact, we can more generally describe the $\mathcal{V}$-linear colimit-preserving approximation of any $\mathcal{V}$-functor. To do this, we first recall the following:
@@
From this perspective, on a given $f$, we can realize this restriction-followed-by-left-Kan-extension as $p_{\mathcal{N}}\circ\mathcal{P}_{\mathcal{V}}(\tilde{f})\circ y$ indeed this functor is $\kappa$-filtered colimit preserving, and agrees with $\tilde{f}$ on $\kappa$-compacts.
Thus we are left with examining the adjunction $\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{M},\mathcal{N})\rightleftarrows\operatorname{Fun}^{\kappa-filt}_{\mathcal{V}}(\mathcal{M},\mathcal{N})$. Using again the description of the latter in terms of $\mathcal{M}^{\kappa}$, and using 1.1, we find that the inclusion is identified with restriction along $p:\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})\to\mathcal{M}$:
@@
${\mathrm{at}}_{\mathcal{M}}(x,-)\simeq\hom_{\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})}(y(x),\hat{y}(-))$
@@
We specialize 2.11 to $\mathcal{N}=\mathcal{V}$. Let $\lambda$ be a cardinal such that $\hom_{\mathcal{M}}(x,-)$ restricted to $\mathcal{M}^{\kappa}$ lands in $\mathcal{V}^{\lambda}$.
We have ${\mathrm{at}}_{\mathcal{M}}(x,-)\simeq p\circ\mathcal{P}_{\mathcal{V}}(\hom_{\mathcal{M}}(x,-))\circ\hat{y}_{\mathcal{M}}$.
@@
Example 2.14.
Let $\mathcal{V}=\mathbf{Sp}$. In this case, we find that ${\mathrm{at}}_{\mathcal{M}}(x,-)\simeq\operatorname{map}(y(x),\hat{y}(-))$. We could call this the spectrum of compact maps, or more generally, in a compactly assembled category, $\operatorname{Map}(y(x),\hat{y}(-))$ could be called the space of compacts maps. Note that its map to $\operatorname{Map}(x,-)$ is not an inclusion of components.
@@
Let $\mathcal{M}$ be a $\mathcal{V}$-module and $x\in\mathcal{M}$. $x$ is atomic if and only if the identity $\mathrm{id}_{x}$ is atomic, if and only the map ${\mathrm{at}}_{\mathcal{M}}(x,-)\to\hom_{\mathcal{M}}(x,-)$ is an equivalence.
@@
It is natural to expect these maps to fit into a more coherent collection of maps and transformations, forming something like an “ideal” in the $\mathcal{V}$-category $\mathcal{M}$. Setting up this amount of coherence would take us too far afield, and we have not had the need for this amount of structure so far. It would certainly be interesting to have it available, to clarifiy the connection between dualizability and ideals (cf.[KNP]).
@@
The universal property from the proof of Example 2.20 together with the equivalence $\hom_{\mathcal{M}}(-,-)\simeq\hom_{\mathcal{N}}(f(-),f(-))$ together give the desired equivalence. ∎
@@
However, over a general base $\mathcal{V}$, two problems arise: first, one cannot simply write objects of $\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})$ as colimits of objects in the image of $y$ one also needs tensors by elements of $\mathcal{V}$, i.e. one essentially needs weighted colimits.
@@
Let $\alpha:x=\mathrm{colim}_{I}^{W}f\to y$ be an atomic presentation of a map in a $\mathcal{V}$-module $\mathcal{M}$, with a chosen lift $\tilde{\alpha}:W\to{\mathrm{at}}_{\mathcal{M}}(f(-),y)$ of the corresponding map of weights $\alpha:W\to\hom_{\mathcal{M}}(f(-),y)$. Let $j:\mathcal{M}\to\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})$ be the restricted Yoneda embedding with left adjoint $p$, where $\kappa$ is an $\mathcal{M}$-good cardinal such that $f$ lands in $\mathcal{M}^{\kappa}$. Finally let $X:=\mathrm{colim}_{I}^{W}j\circ f$. Note that we have a map $X\to j(x)$ given by 1.6.
@@
where the second map uses Example 2.16. Second, we use 2.22 to get [^7] This is where we need $f$ to land in $\mathcal{M}^{\kappa}$., in total, a map
@@
(0 implies 1): Recall that essentially by definition, the canonical functor $p:\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})\to~{}\mathcal{M}$ is given as $\mathrm{colim}_{\mathcal{M}^{\kappa}}^{(-)}i$, where $i:\mathcal{M}^{\kappa}\to\mathcal{M}$ is the inclusion.
Now the map $\hat{y}\to y$ induces equivalences $p\circ\hat{y}\to p\circ y\to\mathrm{id}_{\mathcal{M}}$, and by 2.13 it corresponds exactly to the map ${\mathrm{at}}_{\mathcal{M}}(-,m)\to\hom_{\mathcal{M}}(-,m)$, so the resulting expression
@@
Let $X_{n}:=~{}\mathrm{colim}_{I_{n}}^{W_{n}}y(x^{n}_{\bullet})$. By 2.22, we get maps
@@
To prove this claim, we now use the usual trick: using 2.27, we find dotted lifts in the diagrams:
@@
By 2.28, the corresponding map $\hom(X,\mathrm{colim}_{J}^{V}j\circ g)\to\hom(x,\mathrm{colim}_{J}^{V}g)$ is an equivalence, so the claim is immediate by 2-out-of-3. ∎
@@
Let $V:J^{\mathrm{op}}\to\mathcal{V}$ be a weight and $g:J\to\mathcal{N}$ a diagram. We wish to prove (cf. 1.9) that the canonical map $\mathrm{colim}^{V}_{J}f^{R}\circ g\to f^{R}(\mathrm{colim}^{V}_{J}g)$ is an equivalence. It suffices to check this after applying $\hom(x,-)$ for every $x\in\mathcal{M}$. By 2.25, these are all atomically presentable. So let $x=\mathrm{colim}_{I}^{W}f$ be an atomic presentation of some $x\in\mathcal{M}$ and let $X=\mathrm{colim}_{I}^{W}j\circ f$.
By 2.30 it suffices to prove that the canonical map $\hom(X,\mathrm{colim}_{J}^{V}j\circ f^{R}\circ~{}g)\to~{}\hom(x,f^{R}(\mathrm{colim}_{J}^{V}g))$ is an equivalence.
@@
If $x$ is atomically presentable, 2.28 shows that it is shape-able: if $x\simeq\mathrm{colim}_{I}^{W}f$ is an atomic presentation, we see that $\hat{j}(x)=\mathrm{colim}_{I}^{W}j\circ f$ is the shape of $x$.
Conversely, it is not entirely clear to the author whether shapes are always atomic presentations, namely, whether $\hom_{\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})}(y(-),\hat{y}(x))\to\hom_{\mathcal{M}}(-,x)$ lifts to ${\mathrm{at}}_{\mathcal{M}}(-,x)$: by 2.13, this is true when $\mathcal{M}$ is dualizable, but if not, the fact that $\hat{y}(-)$ is not globally defined makes it difficult to answer this question.
@@
By [Ram24, Proposition 2.31], 2.39 and by construction, $\mathcal{M}^{(\alpha)}$ is $\kappa$-compactly generated for all $\alpha$ and furthermore $\mathcal{M}^{(\alpha)}\to~{}\mathcal{M}$ preserves $\kappa$-compacts.
@@
This is essentially the same proof as in 2.30 and 2.29, so we leave it as an instructive exercise to the reader.
@@
Conversely, 2.46 shows that if $\mathcal{N}\to\mathcal{M}_{\mathrm{dbl}}$ is an internal left adjoint, so is $\mathcal{N}\to\mathcal{M}$. In total, this proves that $\operatorname{Fun}^{iL}_{\mathcal{V}}(\mathcal{N},\mathcal{M}_{\mathrm{dbl}})\to\operatorname{Fun}^{iL}_{\mathcal{V}}(\mathcal{N},\mathcal{M})$ is an equivalence. ∎

### v2

@@ Abstract
We develop the theory of locally rigid and rigid symmetric monoidal $\infty$-categories over an arbitrary base $\mathcal{V}\in\mathrm{CAlg}(\mathbf{Pr}^{\mathrm{L}})$. Among other things, we prove that every locally rigid commutative $\mathcal{V}$-algebra arises as a “completion” of a rigid commutative $\mathcal{V}$-algebra.
Along the way, we introduce and study $\mathcal{V}$-atomic morphisms”, which are analogues of compact morphisms over an arbitrary base $\mathcal{V}$.
@@ Introduction
Typically, one studies not only compactly generated $\infty$-categories, but also (symmetric) monoidal versions of them. Among those, the really well behaved ones are the compactly rigidly generated $\infty$-categories, namely those for which compact objects coincide with dualizable objects. Typical examples include the $\infty$-category of spectra, or more generally the $\infty$-category of $R$-module spectra for any commutative ring spectrum $R$, but also slightly more exotic objects such as the genuine equivariant stable homotopy category $\mathbf{Sp}_{G}$. They are the main objects of study of classical “tt-geometry”, up to the difference between tensor-triangulated categories and stably symmetric monoidal $\infty$-categories.
The rise of dualizable $\infty$-categories therefore requires the rise of a notion of symmetric monoidal dualizable category analogous to this one, without requiring a sufficient supply of compact/dualizable objects. This notion is that of a rigid $\infty$-category, as introduced by Gaitsgory and Rozenblyum, see e.g. [GR19b, Chapter 1, §9] or [HSS+21, Section 2.2] for accounts. We view rigid $\infty$-categories as presentably symmetric monoidal $\infty$-categories that are as close as they could be to being “rigidly-compactly generated” without quite being compactly generated. A good example to keep in mind is the $\infty$-category $\mathbf{Sh}(X)$ of sheaves of spectra on a compact Hausdorff space $X$.
The main goal of this paper is to survey in some depth the foundations of the notion of rigid $\infty$-categories, and how they relate to dualizable $\infty$-categories, and to develop their theory relative to an arbitrary base. As in [RAM24] [^1] Of which the present paper can be viewed as a multiplicative companion, our first result is a presentability result:
@@
Along the way, we are naturally led to the study of atomic morphisms, which are to atomic objects what compact morphisms are to compact objects - or maybe they are to compact morphisms what atomic objects are to compact objects (see [RAM24] for the relevant notions of atomic objects and compact morphisms). In any case, they are a relative notion of compact morphisms. A second goal of this paper is also to exposit the basic properties of atomic morphisms, and try to explain how one might manipulate them - they are slightly more finnicky than compact morphisms in the stable setting, though they have the advantage of working in large generality. To illustrate the analogy, we prove the following theorem, which is meant to generalize [RAM24, Theorem 2.39] - see Section˜2 for precise definitions:
@@
The precise goal of this paper is therefore to study the theory of locally rigid $\infty$-categories, and relate it to that of rigid $\infty$-categories and atomic morphisms.
Our main result in the realm of locally rigid $\infty$-categories (and, arguably, in this paper) is the following result which precisely relates locally rigid and rigid $\infty$-categories- it can be viewed as stating that locally rigid $\infty$-categories all arise as “completions” of rigid ones:
@@
In Section˜1, we review preliminaries concerning weighted colimits in enriched $\infty$-categories, which will be useful later on to set up the notion of “atomic presentation”, which is our relative analogue of the compact exhaustions of [RAM24, Section 2.2].
In Section˜2, we dive into the heart of the matter by introducing atomic morphisms and studying their basic properties, which are meant to be analogous to those of compact morphisms. This is where we prove ˜2.25, namely ˜B from the introduction.
In Section˜3, we recall the definition and basic properties of trace-class morphisms which will be crucial to the study of rigidity.
Section˜4 is the main section of the paper, this is where we introduce and study local rigidity and rigidity. We extend and give new proofs of several results from [AGK+20, Appendix C], and give an analogue of ˜2.25 for rigid categories, namely ˜4.98. Throughout this section, we explain how the story simplifies if we fix $\mathcal{V}=\mathbf{Sp}$, and prove the relevant analogous results. Finally, it is in this section that we prove the existence of, and give a “concrete” construction of rigidifications, allowing us to prove ˜4.67, namely ˜C from the introduction.
Finally, in Section˜5, we do some set theory and prove ˜5.4, namely ˜A.
@@
As mentioned in the introduction, the notion of rigid $\infty$-category comes from the work of Gaitsgory and Rozenblyum [GR19b], and was further studied in [HSS+21]; and that of locally rigid $\infty$-category comes from [AGK+20]. Some of the results we prove here can be found there, at least in the special case $\mathcal{V}$ is itself rigid over $\mathbf{Sp}$.
Some of it is also inspired by the work of Clausen and Scholze on nuclear categories, and some of it also comes directly from the horse’s mouth, such as Section˜4.2 and the corresponding Section˜4.3.1, that I learned directly from Dustin Clausen.
@@ Acknowledgements
The acknowledgements for this paper are essentially the same as for the companion [RAM24]. I specifically need to thank Robert Burklund for asking specifically about locally rigid categories rather than restricting to rigid ones. I furthermore wish to thank Jiacheng Liang for his many helpful comments, and particularly for spotting the mistake recorded in Section˜5.2.
@@ 1 Preliminaries: weighted colimits
where the first equivalence comes from ˜1.1.
@@
This allows to define weighted colimits more generally, as objects satisfying this equivalence if they exist [^5] This is beyond the scope of this paper, but this allows to also define “absolute weights”, which are weights $W$ such that $\mathrm{colim}_{I}^{W}(-)$ is preserved whenever it exists. Once this notion is set up, one may try to compare Karoubi complete $\mathcal{V}$-categories, that is, those $\mathcal{V}$-categories that admit all absolute weighted colimits, and atomically generated $\mathcal{V}$-modules. They should be equivalent along the $\mathcal{P}_{\mathcal{V}}\dashv(-)^{\mathrm{at}}$ adjunction from [BEN23]..
@@
Let $\underline{\mathbf{1}}:(\mathbf{1}_{\mathcal{V}}\otimes I)^{\mathrm{op}}\to\mathcal{V}$ be the constant weight with value $\mathbf{1}_{\mathcal{V}}$. In this case, combining universal properties we find that $\mathrm{colim}_{\mathbf{1}_{\mathcal{V}}\otimes I}^{\underline{\mathbf{1}}}(-):\operatorname{Fun}(I,\mathcal{M})\simeq\operatorname{Fun}_{\mathcal{V}}(\mathbf{1}_{\mathcal{V}}\otimes I,\mathcal{M})\to\mathcal{M}$ is the ordinary $I$-indexed colimit functor.
@@
If $f$ is strong symmetric monoidal and has a right adjoint $^{R}, f#f_{\#}$ is left adjoint to $(f^{R})_{\#}$, and we let $f^{*}$ denote the latter.
@@
One can also give a proof at the level of categories that the resulting functors are $\mathcal{V}$-naturally equivalent.
@@ 2 Atomic maps
In particular, when $\mathcal{V}=\mathbf{Sp}$, we will see that $\mathbf{Sp}$-atomic maps are closely related to compact maps (see Section˜2 for a precise statement).
@@
Proof.
A priori, strongly compact morphisms are defined in terms of filtered colimit preserving functors $f:\mathcal{M}\to\mathbf{S}$ and maps $f\to\operatorname{Map}(x,-)$, while $\mathbf{Sp}$-atomic morphisms are defined in terms of filtered colimit preserving, exact functors $g:\mathcal{M}\to\mathbf{Sp}$ and maps $g\to\operatorname{map}(x,-)$. Thus, it is clear that a $\mathbf{Sp}$-atomic map is strongly compact, since $\Omega^{\infty}$ preserves filtered colimits and sends $\operatorname{map}(x,-)$ to $\operatorname{Map}(x,-)$.
On the other hand, $\operatorname{Map}(x,-):\mathcal{M}\to\mathbf{S}$ is a $1$-excisive reduced functor, so if $f:\mathcal{M}\to\mathbf{S}$ is filtered colimit preserving and has a map $f\to\operatorname{Map}(x,-)$, then taking an excisive approximation of $f$ and using the fact that exact functors $\mathcal{M}\to\mathbf{Sp}$ are equivalent to excisive reduced functors $\mathcal{M}\to\mathbf{S}$, we find that $f\to\operatorname{Map}(x,-)$ factors through a map obtained as $\Omega^{\infty}$ of a map $g\to\operatorname{map}(x,-)$, where $g:\mathcal{M}\to\mathbf{Sp}$ is filtered colimit preserving [^6] This uses that excisive approximations of filtered colimit-preserving functors preserve filtered colimits. and exact. Thus, a strongly compact map is $\mathbf{Sp}$-atomic. ∎
@@
The first half of the claim is Section˜2. The second half follows from the first together with [RAM24, Example 2.20]. ∎
@@
As a consequence, ${\mathrm{at}}_{\mathcal{V}}(x,-)\simeq\hom_{\mathcal{V}}(x,\mathbf{1}_{\mathcal{V}})\otimes$-, and so $\mathcal{V}$-atomic maps in $\mathcal{V}$ are the same as trace-class maps (see Section˜3 and Section˜3).
@@
We can describe ${\mathrm{at}}_{\mathcal{M}}(x,-)$ a bit more explicitly when $\mathcal{M}$ is dualizable over $\mathcal{V}$ - in fact, we can more generally describe the $\mathcal{V}$-linear colimit-preserving approximation of any $\mathcal{V}$-functor. To do this, we first recall the following:
@@
From this perspective, on a given $f$, we can realize this restriction-followed-by-left-Kan-extension as $p_{\mathcal{N}}\circ\mathcal{P}_{\mathcal{V}}(\tilde{f})\circ y$ - indeed this functor is $\kappa$-filtered colimit preserving, and agrees with $\tilde{f}$ on $\kappa$-compacts.
Thus we are left with examining the adjunction $\operatorname{Fun}^{L}_{\mathcal{V}}(\mathcal{M},\mathcal{N})\rightleftarrows\operatorname{Fun}^{\kappa-filt}_{\mathcal{V}}(\mathcal{M},\mathcal{N})$. Using again the description of the latter in terms of $\mathcal{M}^{\kappa}$, and using ˜1.1, we find that the inclusion is identified with restriction along $p:\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})\to\mathcal{M}$:
@@
[…]
@@
We specialize Section˜2 to $\mathcal{N}=\mathcal{V}$. Let $\lambda$ be a cardinal such that $\hom_{\mathcal{M}}(x,-)$ restricted to $\mathcal{M}^{\kappa}$ lands in $\mathcal{V}^{\lambda}$.
We have $^{\mathrm{at}}_{\mathcal{M}}(x,-)\simeq p\circ\mathcal{P}_{\mathcal{V}}(\hom_{\mathcal{M}}(x,-))\circ\hat{y}_{\mathcal{M}}$.
But now, $\hom_{\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})}(y(x),-)$ and $p\circ\mathcal{P}_{\mathcal{V}}(\hom_{\mathcal{M}}(x,-))$ by design agree on $\mathcal{M}^{\kappa}\xhookrightarrow{y}~\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})$ and are both $\mathcal{V}$-linear and colimit-preserving [^7] This is where we use that $x$ is in $\mathcal{M}^{\kappa}$: otherwise $\hom(y(x),-)$ need not preserve colimits or be $\mathcal{V}$-linear.. Thus they agree, and we are done. ∎
Example 2.14.
@@
Let $\mathcal{M}$ be a $\mathcal{V}$-module and $\in\mathcal{M}. xx$ is atomic if and only if the identity $\mathrm{id}_{x}$ is atomic, if and only the map ${\mathrm{at}}_{\mathcal{M}}(x,-)\to\hom_{\mathcal{M}}(x,-)$ is an equivalence.
@@
It is natural to expect these maps to fit into a more coherent collection of maps and transformations, forming something like an “ideal” in the $\mathcal{V}$-category $\mathcal{M}$. Setting up this amount of coherence would take us too far afield, and we have not had the need for this amount of structure so far. It would certainly be interesting to have it available, to clarifiy the connection between dualizability and ideals (cf.[11]).
@@
The universal property from the proof of Section˜2 together with the equivalence $\hom_{\mathcal{M}}(-,-)\simeq\hom_{\mathcal{N}}(f(-),f(-))$ together give the desired equivalence. ∎
@@
However, over a general base $\mathcal{V}$, two problems arise: first, one cannot simply write objects of $\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})$ as colimits of objects in the image of $y$ - one also needs tensors by elements of $\mathcal{V}$, i.e. one essentially needs weighted colimits.
@@
Let $\alpha:x=\mathrm{colim}_{I}^{W}f\to y$ be an atomic presentation of a map in a $\mathcal{V}$-module $\mathcal{M}$, with a chosen lift $\tilde{\alpha}:W\to{\mathrm{at}}_{\mathcal{M}}(f(-),y)$ of the corresponding map of weights $\alpha:W\to\hom_{\mathcal{M}}(f(-),y)$. Let $j:\mathcal{M}\to\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})$ be the restricted Yoneda embedding with left adjoint $p$, where $\kappa$ is an $\mathcal{M}$-good cardinal such that $f$ lands in $\mathcal{M}^{\kappa}$. Finally let $X:=\mathrm{colim}_{I}^{W}j\circ f$. Note that we have a map $X\to j(x)$ given by Section˜1.
@@
where the second map uses Section˜2. Second, we use Section˜2 to get [^8] This is where we need $f$ to land in $\mathcal{M}^{\kappa}$., in total, a map
@@
(0 implies 1): Recall that essentially by definition, the canonical functor $p:\mathcal{P}_{\mathcal{V}}(\mathcal{M}^{\kappa})\to$$\mathcal{M}$ is given as $\mathrm{colim}_{\mathcal{M}^{\kappa}}^{(-)}i$, where $i:\mathcal{M}^{\kappa}\to\mathcal{M}$ is the inclusion.
Now the map $\hat{y}\to y$ induces equivalences $p\circ\hat{y}\to p\circ y\to\mathrm{id}_{\mathcal{M}}$, and by Section˜2 it corresponds exactly to the map ${\mathrm{at}}_{\mathcal{M}}(-,m)\to\hom_{\mathcal{M}}(-,m)$, so the resulting expression
@@
Let $X_{n}:=~\mathrm{colim}_{I_{n}}^{W_{n}}y(x^{n}_{\bullet})$. By Section˜2, we get maps
@@
To prove this claim, we now use the usual trick: using Section˜2, we find dotted lifts in the diagrams:
@@
By Section˜2, the corresponding map $\hom(X,\mathrm{colim}_{J}^{V}j\circ g)\to\hom(x,\mathrm{colim}_{J}^{V}g)$ is an equivalence, so the claim is immediate by 2-out-of-3. ∎
@@
Let $V:J^{\mathrm{op}}\to\mathcal{V}$ be a weight and $g:J\to\mathcal{N}$ a diagram. We wish to prove (cf. Section˜1) that the canonical map $\mathrm{colim}^{V}_{J}f^{R}\circ g\to f^{R}(\mathrm{colim}^{V}_{J}g)$ is an equivalence. It suffices to check this after applying $\hom(x,-)$ for every $x\in\mathcal{M}$. By ˜2.25, these are all atomically presentable. So let $x=\mathrm{colim}_{I}^{W}f$ be an atomic presentation of some $x\in\mathcal{M}$ and let $X=\mathrm{colim}_{I}^{W}j\circ f$.
By Section˜2 it suffices to prove that the canonical map $\hom(X,\mathrm{colim}_{J}^{V}j\circ f^{R}\circ~g)\to~\hom(x,f^{R}(\mathrm{colim}_{J}^{V}g))$ is an equivalence.
@@
If $x$ is atomically presentable, Section˜2 shows that it is shape-able: if $x\simeq\mathrm{colim}_{I}^{W}f$ is an atomic presentation, we see that $\hat{j}(x)=\mathrm{colim}_{I}^{W}j\circ f$ is the shape of $x$.
[…]
@@
By [RAM24, Proposition 2.31], Section˜2 and by construction, $\mathcal{M}^{(\alpha)}$ is $\kappa$-compactly generated for all $\alpha$ and furthermore $\mathcal{M}^{(\alpha)}\to$$\mathcal{M}$ preserves $\kappa$-compacts.
@@
This is essentially the same proof as in Section˜2 and Section˜2, so we leave it as an instructive exercise to the reader.
@@
Conversely, Section˜2 shows that if $\mathcal{N}\to\mathcal{M}_{\mathrm{dbl}}$ is an internal left adjoint, so is $\mathcal{N}\to\mathcal{M}$. In total, this proves that $\operatorname{Fun}^{iL}_{\mathcal{V}}(\mathcal{N},\mathcal{M}_{\mathrm{dbl}})\to\operatorname{Fun}^{iL}_{\mathcal{V}}(\mathcal{N},\mathcal{M})$ is an equivalence. ∎